# A Coherence-Based Measure of AGI

### v1

@@ 3 A Coherence-Based Measure of AGI
To summarize model performance across compensability regimes, we define the AUC-based AGI score:
$\mathrm{AGI}_{\mathrm{AUC}}=\frac{1}{p_{\max}-p_{\min}}\int_{p_{\min}}^{p_{\max}}\mathrm{AGI}_{p}\,dp,$ (3)
where $\mathrm{AGI}_{p}$ is evaluated for $p\in[p_{\min},p_{\max}]$. In this work, we integrate over the interval [1,1][-1,1], spanning from the harmonic mean to the arithmetic mean. $\mathrm{AGI}_{\mathrm{AUC}}$ quantifies the average coherence of intelligence across coupling regimes: higher values indicate robustness under stricter (more non-compensatory) conditions, whereas lower values reveal dependence on a small subset of specialized abilities.
The integral was computed numerically using the composite trapezoidal rule over a uniform grid of $p\in[-1,1]$, providing a stable and reproducible approximation. This approach ensures smooth sensitivity to local variations in $\mathrm{AGI}_{p}$ and avoids bias toward any particular regime.
Current AI systems often exhibit highly uneven or “jagged” ability profiles (Hendrycks et al., 2025), performing near human level in some domains while failing catastrophically in others. The $\mathrm{AGI}_{p}$ family makes this imbalance explicit: $\mathrm{AGI}_{1}$ captures the breadth of capability, whereas $\mathrm{AGI}_{\mathrm{AUC}}$ summarizes the stability of capability across compensability regimes—offering a stricter, coherence-sensitive indicator of general intelligence.
4 Results
@@
We apply the proposed framework to the published AGI domain scores of Hendrycks et al. (2025) for GPT-4 (Achiam et al., 2023) and GPT-5 (2025). For completeness, these baseline scores are reproduced in Table 1. The floor constant was fixed at $\varepsilon=10^{-6}$, and the compensability interval was set to $p\in[-1,1]$. We report $\mathrm{AGI}_{-1}$, $\mathrm{AGI}_{-0.5}$, $\mathrm{AGI}_{0}$, $\mathrm{AGI}_{0.5}$, and $\mathrm{AGI}_{1}$, along with their corresponding $\mathrm{AGI}_{p}$ curves and the aggregate $\mathrm{AGI}_{\mathrm{AUC}}$ measure. To assess construct validity, we further compare the internal $\mathrm{AGI}_{\mathrm{AUC}}$ metric against challenging external reasoning benchmarks.
Our main analysis builds directly on the ten normalized domain scores reported in Hendrycks et al. (2025), which serve as the empirical foundation for this study. In Appendix A, we document the origin of these values and compute several alternative aggregates, including weighted and unweighted geometric means. We keep this extended discussion in the appendix to maintain focus on the primary contribution of this paper: demonstrating how coherence-aware aggregation provides a more rigorous and interpretable framework for assessing general intelligence.
Figure 1: Comparison of model performance across aggregation exponents pp. Curves show $\mathrm{AGI}_{p}$ values; shaded regions indicate AUC-based uncertainty envelopes.
@@ Results.
Figure 1 and Table 2 summarize how model performance varies under different compensability assumptions. When $p=1$, the arithmetic mean ($\mathrm{AGI}_{1}$) reproduces the values reported in Hendrycks et al. (2025). Under this compensatory definition, GPT-5 appears to have made substantial progress over GPT-4 (58% vs. 27%), suggesting meaningful gains in overall capability and surpassing the halfway mark toward human-level proficiency.
However, as compensability decreases ($p\leq 0$), this apparent advantage diminishes sharply. The geometric mean ($\mathrm{AGI}_{0}$) collapses to near zero for both GPT-4 and GPT-5, reflecting the persistence of minimal proficiency in several cognitive domains. This pattern indicates that both systems rely heavily on a subset of strong faculties rather than demonstrating balanced competence across domains. Consequently, the arithmetic mean tends to overstate progress by averaging across highly uneven abilities, producing an inflated signal of movement toward AGI.
The area-under-the-curve measure, $\mathrm{AGI}_{\mathrm{AUC}}$, integrates performance across $p\in[-1,1]$ to capture overall coherence. GPT-5 achieves a roughly threefold improvement over GPT-4 (24% vs. 7%), reflecting greater but still limited robustness. The $\mathrm{AGI}_{p}$ curves in Figure 1 exhibit steep declines for negative $p$ values, indicating that weaknesses in domains such as long-term memory and adaptive reasoning continue to act as system-wide bottlenecks. A system approaching genuine generality would instead maintain a flatter and higher $\mathrm{AGI}_{p}$ curve across the full range of $p$, indicating resilience under increasingly strict coupling regimes. In the speculative case of “GPT-6?”, shown in In Figure 1, we simulate the effect of mitigating GPT-5’s principal bottleneck in long-term memory storage (MS; see Table 1) by increasing its domain score from 0% to 30%, while keeping all other domain scores unchanged. Both $\mathrm{AGI}_{p}$ and $\mathrm{AGI}_{\mathrm{AUC}}$ increase substantially, suggesting that even modest improvements in the weakest faculties can yield disproportionate gains in overall coherence as captured by the proposed metrics. This illustrative scenario highlights the non-compensatory structure of general intelligence: closing a single critical deficit can markedly enhance systemic robustness across all coupling regimes.
Alignment with external benchmarks.
To assess construct validity, we compared the internal $\mathrm{AGI}_{p}$ metrics with challenging external reasoning benchmarks. The ARC-AGI-2 benchmark (Chollet, 2019; Chollet et al., 2025), which emphasizes out-of-distribution reasoning and cross-domain abstraction, reveals a clear calibration effect. While GPT-5 achieves an arithmetic-mean score of $\mathrm{AGI}_{1}={\color[rgb]{0,0.4,0}58}\%$, suggesting near-human generality, its coherence-aware score of $\mathrm{AGI}_{\mathrm{AUC}}={\color[rgb]{0.4,0,0}24}\%$ aligns much more closely with ARC-AGI-2 results, where GPT-5 (Pro) reaches approximately 18% and GPT-5 (High) about 10%. Similarly, the BIG-Bench Extra Hard evaluations (Kazemi et al., 2025), which probe broader reasoning proficiency across linguistically diverse tasks, report a GPT-4 score of roughly 6%—closely matching its $\mathrm{AGI}_{\mathrm{AUC}}={\color[rgb]{0.4,0,0}7}\%$, unlike the arithmetic mean of 27%.
This correspondence supports the interpretation that $\mathrm{AGI}_{\mathrm{AUC}}$ more faithfully captures functional coherence—the capacity to sustain competence across diverse cognitive domains—whereas the arithmetic mean $\mathrm{AGI}_{1}$ tends to overstate progress by rewarding isolated peaks of specialization. Accordingly, $\mathrm{AGI}_{\mathrm{AUC}}$ serves as a more construct-valid and conservative indicator of emerging generality, aligning closely with independent measures of out-of-distribution reasoning such as ARC-AGI-2 and BIG-Bench Hard.
@@ 5 Discussion
The results indicate that the arithmetic mean used in prior definitions of AGI ($\mathrm{AGI}_{1}$) is a fragile indicator of generality. It rewards specialization and conceals structural weaknesses. Under coherence-aware formulations ($\mathrm{AGI}_{0}$, $\mathrm{AGI}_{-1}$, and $\mathrm{AGI}_{\text{AUC}}$), the same systems score far lower, revealing that contemporary models remain unevenly developed across core cognitive domains.
The current score is stricter yet still optimistic.
The coherence-aware framework reveals that current estimates of general intelligence in frontier models are likely inflated by additive aggregation. While GPT-5 achieves an $\mathrm{AGI}_{1}$ (arithmetic mean) of $\color[rgb]{0,0.4,0}58$$\%$, suggesting near-human breadth, its coherence-adjusted $\mathrm{AGI}_{\mathrm{AUC}}$ falls to only $\color[rgb]{0.4,0,0}24$$\%$. Even this value should be regarded as an upper bound on general competence. As shown in Appendix A, the ten constituent cognitive scores from Hendrycks et al. (2025) themselves display substantial inflation when aggregated arithmetically. When recalculated using geometric or unweighted means, several domains collapse: auditory processing drops from roughly $\color[rgb]{0,0.4,0}60$$\%$ to near 0%, visual processing from $\color[rgb]{0.4,0,0}40$$\%$ to near $\color[rgb]{0.4,0,0}0$$\%$, processing speed from $\color[rgb]{0.4,0,0}30$$\%$ to near $\color[rgb]{0.4,0,0}0$$\%$, and on-the-spot reasoning from $\color[rgb]{0,0.4,0}70$$\%$ (arithmetic) to $\color[rgb]{0.4,0,0}19$$\%$ (weighted geometric) or $\color[rgb]{0.4,0,0}5$$\%$ (unweighted). These discrepancies expose a systematic overestimation of generality due to compensatory aggregation, where strong subskills mask critical deficiencies elsewhere. Accordingly, the nominal $\mathrm{AGI}_{\mathrm{AUC}}$ of $\color[rgb]{0.4,0,0}24$$\%$ should be interpreted as conservative yet still optimistic, given (potential) inflation at the domain-score level.
@@
The divergence between $\mathrm{AGI}_{1}$ and coherence-based metrics highlights a key distinction between breadth of performance and integration of capability. Improvements in select domains inflate arithmetic averages but contribute little to genuine generality if other faculties remain weak. This mirrors refinements in human psychometrics, where single-factor IQ models evolved into hierarchical frameworks (Carroll, 1993; McGrew, 2009, 2023a) to capture both domain-specific and global ability. AI evaluation requires a similar shift—from measuring average proficiency to assessing functional sufficiency across interdependent cognitive dimensions.
@@
The generalized-mean formulation offers a continuous lens for exploring compensability assumptions. Across this spectrum, $\mathrm{AGI}_{\text{AUC}}$ emerges as the most stable and informative single measure: it integrates performance across optimistic and strict regimes, balancing sensitivity to weaknesses with tolerance for partial unevenness. Systems that achieve high $\mathrm{AGI}_{1}$ but low $\mathrm{AGI}_{0}$ or $\mathrm{AGI}_{\text{AUC}}$ should thus be interpreted as specialized rather than coherent generalists. Conversely, genuine progress toward AGI would manifest as a flatter, consistently high $\mathrm{AGI}_{p}$ curve— indicating resilience under increasingly strict coupling of abilities. Future benchmarks and model cards should therefore report both $\mathrm{AGI}_{1}$ (average proficiency) and coherence-oriented measures such as $\mathrm{AGI}_{\text{AUC}}$ to disentangle breadth from balance.
@@
Although coherence-based measures expose the structural limitations of arithmetic averaging, intelligence cannot be reduced to any single scalar quantity. Each aggregate reflects a complementary facet of generality: $\mathrm{AGI}_{1}$ captures breadth, $\mathrm{AGI}_{0}$ reflects balance, and $\mathrm{AGI}_{\text{AUC}}$ summarizes stability across compensability regimes. Together, they form a multidimensional diagnostic toolkit rather than a single leaderboard number. External benchmarks such as ARC-AGI-2 (Chollet et al., 2025) and BIG-Bench Extra Hard (Kazemi et al., 2025), among several others, further complement these CHC-based metrics by probing abstraction and reasoning beyond training distributions.
@@
Low geometric or AUC-based scores do not imply an absence of generalization, but rather highlight fragility when tasks demand multi-domain coordination. The coherence perspective thus bridges psychometric and computational traditions, emphasizing integration and stability over isolated domain success. The goal of proposing a stricter, coherence-based measure is not to diminish the remarkable capabilities of current GPT-class systems, but to sharpen the lens through which progress toward AGI is assessed. True general intelligence should demonstrate balanced competence in reasoning, abstraction, generalization, planning, and memory—the capacities that underpin adaptability and understanding beyond pattern replication. By tightening the definition of generality, this framework aims to steer progress toward integrated, cross-domain coherence—an intelligence that not only performs well, but reasons, learns, and plans as a unified system.
@@
The proposed measure inherits a key strength of the prior definition Hendrycks et al. (2025): it evaluates systems at the task rather than the dataset level. Our operationalization centers on task specifications, with datasets serving only as illustrative—and thus necessary but not sufficient—instances of those tasks. Solving a benchmark does not entail solving the underlying task, since automatic evaluations often underrepresent their target phenomena Yogatama et al. (2019). By grounding evaluation in task-level constructs, the framework remains robust and adaptable, accommodating new or improved benchmarks without altering its underlying definition of general intelligence.
Limitations and future work.
Our formulation of $\mathrm{AGI}_{\mathrm{AUC}}$ builds directly on the ten normalized domain scores reported in Hendrycks et al. (2025), which serve as the empirical foundation for this analysis. These values provide a common reference point across domains but also inherit limitations from that taxonomy and weighting methodology. As shown in Appendix A, arithmetic-based scores may be inflated relative to their geometric counterparts, emphasizing the sensitivity of evaluation to aggregation choice. Future work should refine the domain taxonomy, introduce nonlinear scaling and adaptive weightings based on empirical task dependencies, and track $\mathrm{AGI}_{p}$ evolution across model generations. Such efforts could yield diagnostic tools identifying which deficiencies most constrain $\mathrm{AGI}_{\mathrm{AUC}}$, guiding targeted architectural and training interventions to enhance systemic coherence.

### v2

@@ 3.2 An Integrated Measure of Coherence: $\mathrm{AGI}_{\mathrm{AUC}}$
To summarize robustness across compensability regimes, we define the $\mathrm{AGI}_{\mathrm{AUC}}$ metric:
$\mathrm{AGI}_{\mathrm{AUC}}=\frac{1}{p_{\max}-p_{\min}}\int_{p_{\min}}^{p_{\max}}\mathrm{AGI}_{p}\,dp.$ (3)
Throughout this work, we set $p_{\min}=-1$ and $p_{\max}=1$, spanning the range from strongly non-compensatory to strongly compensatory aggregation. The integral can be evaluated numerically using the composite trapezoidal rule over a uniform grid in $p$.
The $\mathrm{AGI}_{p}$ family highlights the characteristic “jaggedness” of modern AI systems: high proficiency in some domains coexists with catastrophic failures in others.
$\mathrm{AGI}_{\mathrm{AUC}}$ measures coherence and stability across compensability regimes. Systems with high $\mathrm{AGI}_{1}$ but low $\mathrm{AGI}_{\mathrm{AUC}}$ rely on narrow excellence and lack balanced, integrated capability. In contrast, a system approaching general intelligence would display a uniformly high $\mathrm{AGI}_{p}$ curve across the entire range of $p$, indicating robustness even under stricter non-compensatory evaluation. Thus, $\mathrm{AGI}_{\mathrm{AUC}}$ provides a principled and coherence-sensitive summary of general intelligence, capturing both the breadth and the structural balance of a system’s abilities.
4 Results on the CHC-Based Domain Scores
@@
We apply the proposed framework to the published AGI domain scores of hendrycks2025agidefinition for GPT-4 (achiam2023gpt) and GPT-5 (2025). For completeness, these baseline scores are reproduced in Table 1. The floor constant was fixed at $\varepsilon=10^{-6}$, and the compensability interval was set to $p\in[-1,1]$. We report $\mathrm{AGI}_{-1}$, $\mathrm{AGI}_{-0.5}$, $\mathrm{AGI}_{0}$, $\mathrm{AGI}_{0.5}$, and $\mathrm{AGI}_{1}$, along with their corresponding $\mathrm{AGI}_{p}$ curves and the aggregate $\mathrm{AGI}_{\mathrm{AUC}}$ measure. To assess external alignment, we also compare the resulting $\mathrm{AGI}_{\mathrm{AUC}}$ values with performance on challenging intelligence benchmarks.
The analysis in this section builds directly on the ten normalized domain scores reported in hendrycks2025agidefinition. In Appendix A, we document the origin of these values and compute several alternative aggregates, including weighted and unweighted geometric means. We keep this extended discussion in the appendix to maintain focus on the primary contribution of this paper: demonstrating how coherence-based aggregation provides a more principled framework for assessing AGI.
@@ Results.
Figure 2 and Table 2 summarize how model performance varies under different compensability assumptions. When $p=1$, the arithmetic mean ($\mathrm{AGI}_{1}$) reproduces the values reported in hendrycks2025agidefinition. Under this compensatory definition, GPT-5 appears to have made substantial progress over GPT-4 (58% vs. 27%), suggesting meaningful gains in overall capability and surpassing the halfway mark toward human-level proficiency.
However, as compensability decreases, this apparent advantage diminishes sharply. The geometric mean ($\mathrm{AGI}_{0}$) collapses to near zero for both GPT-4 and GPT-5, reflecting the persistence of minimal proficiency in several cognitive domains. This pattern indicates that both systems rely heavily on a subset of strong faculties rather than demonstrating balanced competence across domains. Consequently, the arithmetic mean tends to overstate progress by averaging across highly uneven abilities, producing an inflated signal of movement toward AGI.
The area-under-the-curve measure, $\mathrm{AGI}_{\mathrm{AUC}}$, integrates performance across $p\in[-1,1]$ to capture overall coherence. GPT-5 achieves a roughly threefold improvement over GPT-4 (24% vs. 7%), reflecting greater but still limited robustness. The $\mathrm{AGI}_{p}$ curves in Figure 2 exhibit steep declines for negative $p$ values, indicating that weaknesses in domains such as long-term memory and adaptive reasoning continue to act as system-wide bottlenecks. A system approaching genuine generality would instead maintain a flatter and higher $\mathrm{AGI}_{p}$ curve across the full range of $p$, indicating resilience under increasingly strict coupling regimes. In the speculative case of “GPT-6?”, shown in In Figure 2, we simulate the effect of mitigating GPT-5’s principal bottleneck in long-term memory storage (MS; see Table 1) by increasing its domain score from 0% to 30%, while keeping all other domain scores unchanged. Both $\mathrm{AGI}_{p}$ and $\mathrm{AGI}_{\mathrm{AUC}}$ increase substantially, suggesting that even modest improvements in the weakest faculties can yield disproportionate gains in overall coherence as captured by the proposed metrics.
Alignment with external benchmarks.
We compare the $\mathrm{AGI}_{p}$ metrics with challenging external reasoning benchmarks. The ARC-AGI-2 benchmark (chollet2019measure; chollet2025arc), which emphasizes out-of-distribution reasoning and cross-domain abstraction, reveals a clear calibration effect. While GPT-5 achieves an arithmetic-mean score of $\mathrm{AGI}_{1}={\color[rgb]{0,0.4,0}58}\%$, suggesting near-human generality, its coherence-aware score of $\mathrm{AGI}_{\mathrm{AUC}}={\color[rgb]{0.4,0,0}24}\%$ aligns much more closely with ARC-AGI-2 results, where GPT-5 (Pro) reaches approximately 18% and GPT-5 (High) about 10%. Similarly, the BIG-Bench Extra Hard evaluations (kazemi2025big), which probe broader reasoning proficiency across linguistically diverse tasks, report a GPT-4 score of roughly 6%, closely matching its $\mathrm{AGI}_{\mathrm{AUC}}={\color[rgb]{0.4,0,0}7}\%$, unlike the arithmetic mean of 27%.
This correspondence supports the interpretation that $\mathrm{AGI}_{\mathrm{AUC}}$ more faithfully captures functional coherence, the capacity to sustain competence across diverse cognitive domains, whereas the arithmetic mean $\mathrm{AGI}_{1}$ tends to overstate progress by rewarding isolated peaks of specialization. Accordingly, $\mathrm{AGI}_{\mathrm{AUC}}$ serves as a more conservative indicator of emerging generality, aligning closely with independent measures of out-of-distribution reasoning such as ARC-AGI-2 and BIG-Bench Hard.
5 Beyond CHC-Domain Evaluation
While the previous section focuses on the ten CHC-inspired cognitive domains defined by hendrycks2025agidefinition, the coherence-based framework developed in this work is not tied to that taxonomy. As demonstrated in Appendix B, we further evaluate the same generalized-mean family $\mathrm{AGI}_{p}$ and the integrated coherence metric $\mathrm{AGI}_{\mathrm{AUC}}$ on a heterogeneous set of 17 benchmarks drawn from the publicly released Gemini 3 Pro Model Evaluation Report (deepmind2025gemini3pro), for GPT-5.1, Gemini 3 Pro, Claude Sonnet 2.5 Pro, and Claude Sonnet 4.5 (claude_sonnet4_5). These benchmarks span symbolic reasoning, scientific problem solving, mathematical tasks, coding and agentic behavior, multimodal abstraction, and long-context understanding, illustrating the flexibility of the framework across evaluation settings.
The extended analysis reveals patterns that closely mirror those observed in the CHC-domain setting. Aggregation with $p=1$ (i.e., the arithmetic mean) again yields optimistic assessments of overall capability by masking imbalances across tasks, whereas negative-$p$ aggregation exposes persistent bottlenecks. Models with similar average performance diverge sharply in coherence: their $\mathrm{AGI}_{p}$ curves show steep declines for $p<0$, and $\mathrm{AGI}_{\mathrm{AUC}}$ provides a clearer and more stable ordering of holistic competence. These results reinforce that coherence captures whether a system’s abilities form a balanced, general-purpose profile rather than a collection of narrow skill peaks.
At the same time, while coherence-based aggregation provides a substantially more principled and informative summary than the arithmetic mean, any benchmark-derived score must ultimately be interpreted in the context of the benchmark suite it aggregates. To meaningfully quantify “progress toward AGI,” the benchmarks themselves must span the full spectrum of essential cognitive faculties. Expanding benchmark coverage—across continual learning, long-horizon reasoning, memory formation, planning, and other fundamental abilities, remains an important direction for the community. Within any given benchmark set, however, our coherence-based aggregation offers a robust and conservative measure of generality: it penalizes uneven skill profiles, highlights hidden bottlenecks, and rewards the balance and robustness that are core to general intelligence.
@@ 6 Discussion
The results indicate that the arithmetic mean used in prior definitions of AGI ($\mathrm{AGI}_{1}$) is a fragile indicator of generality. It rewards specialization and conceals structural weaknesses. Under coherence-aware formulations ($\mathrm{AGI}_{0}$, $\mathrm{AGI}_{-1}$, and $\mathrm{AGI}_{\text{AUC}}$), the same systems score far lower, revealing that contemporary models remain unevenly developed across core cognitive domains and skills.
The current CHC-derived score is stricter yet still optimistic.
The coherence-based framework reveals that current estimates of general intelligence in frontier models are likely inflated by additive aggregation. While GPT-5 achieves an $\mathrm{AGI}_{1}$ (arithmetic mean) of $\color[rgb]{0,0.4,0}58$$\%$, suggesting near-human breadth, its coherence-adjusted $\mathrm{AGI}_{\mathrm{AUC}}$ falls to only $\color[rgb]{0.4,0,0}24$$\%$. Even this value should be regarded as an upper bound on general competence. As shown in Appendix A, the ten constituent cognitive scores from hendrycks2025agidefinition themselves display substantial inflation when aggregated arithmetically. When recalculated using geometric or unweighted means, several domains collapse: auditory processing drops from roughly $\color[rgb]{0,0.4,0}60$$\%$ to near 0%, visual processing from $\color[rgb]{0.4,0,0}40$$\%$ to near $\color[rgb]{0.4,0,0}0$$\%$, processing speed from $\color[rgb]{0.4,0,0}30$$\%$ to near $\color[rgb]{0.4,0,0}0$$\%$, and on-the-spot reasoning from $\color[rgb]{0,0.4,0}70$$\%$ (arithmetic) to $\color[rgb]{0.4,0,0}19$$\%$ (weighted geometric) or $\color[rgb]{0.4,0,0}5$$\%$ (unweighted). These discrepancies expose a systematic overestimation of generality due to compensatory aggregation, where strong subskills mask critical deficiencies elsewhere. Accordingly, the nominal $\mathrm{AGI}_{\mathrm{AUC}}$ of $\color[rgb]{0.4,0,0}24$$\%$ should be interpreted as conservative yet still optimistic, given (potential) inflation at the domain-score level.
@@
The divergence between $\mathrm{AGI}_{1}$ and coherence-based metrics highlights a key distinction between breadth of performance and integration of capability. Improvements in select domains inflate arithmetic averages but contribute little to genuine generality if other faculties remain weak. This mirrors refinements in human psychometrics, where single-factor IQ models evolved into hierarchical frameworks (carroll1993human; mcgrew2009chc; mcgrew2023carroll) to capture both domain-specific and global ability. AI evaluation requires a similar shift, from measuring average proficiency to assessing functional sufficiency across interdependent cognitive dimensions.
@@
The generalized-mean formulation offers a continuous lens for exploring compensability assumptions. Across this spectrum, $\mathrm{AGI}_{\text{AUC}}$ emerges as the most stable and informative single measure: it integrates performance across optimistic and strict regimes, balancing sensitivity to weaknesses with tolerance for partial unevenness. Systems that achieve high $\mathrm{AGI}_{1}$ but low $\mathrm{AGI}_{0}$ or $\mathrm{AGI}_{\text{AUC}}$ should thus be interpreted as specialized rather than coherent generalists. Conversely, genuine progress toward AGI would manifest as a flatter, consistently high $\mathrm{AGI}_{p}$ curve, indicating resilience under increasingly strict coupling of abilities. Future benchmarks and model cards should therefore report both $\mathrm{AGI}_{1}$ (average proficiency) and coherence-oriented measures such as $\mathrm{AGI}_{\text{AUC}}$ to disentangle breadth from balance.
@@
Although coherence-based measures expose the structural limitations of arithmetic averaging, no single scalar, including $\mathrm{AGI}_{\mathrm{AUC}}$, can fully capture general intelligence. Each aggregate reflects a complementary facet of capability: the arithmetic mean highlights breadth, negative-$p$ regimes reveal bottlenecks, and the full $\mathrm{AGI}_{p}$ curve provides a visual diagnostic of how performance behaves under varying compensability assumptions. Rather than serving as a single leaderboard number, the continuum over $p$ forms a multidimensional assessment toolkit. External benchmarks such as ARC-AGI-2 (chollet2025arc) and BIG-Bench Extra Hard (kazemi2025big) further complement this analysis, and can themselves be integrated using the same generalized-mean framework to yield coherence-aware summaries of performance.
@@
Low curves or AUC-based scores do not imply an absence of generalization, but rather highlight fragility when tasks demand multi-domain coordination. The coherence perspective thus bridges psychometric and computational traditions, emphasizing integration and stability over isolated domain success. The goal of proposing a stricter, coherence-based measure is not to diminish the remarkable capabilities of current GPT-class systems, but to sharpen the lens through which progress toward AGI is assessed. True general intelligence should demonstrate balanced competence in reasoning, abstraction, generalization, planning, and memory, the capacities that underpin adaptability and understanding beyond pattern replication. By tightening the definition of generality, this framework aims to steer progress toward integrated, cross-domain coherence, an intelligence that not only performs well, but reasons, learns, and plans as a unified system.
@@
A central advantage of the proposed coherence-based measure is its generality: it applies uniformly across heterogeneous benchmark families without committing to a fixed cognitive taxonomy. Our formulation allows for any collection of tasks, whether CHC-aligned domains (as in Section 4), symbolic reasoning suites, multimodal assessments, or mixed benchmark portfolios (as in Section 5), to be aggregated using the same generalized-mean family $\mathrm{AGI}_{p}$. The metric depends only on normalized task scores, not on assumptions about human-inspired structure or domain boundaries. As demonstrated in Appendix B, this enables seamless integration of diverse evaluations (e.g., Humanity’s Last Exam (phan2025humanity) and ARC-AGI-2 (chollet2025arc)), producing coherence-based summaries that remain practical even as benchmarks evolve. By grounding assessment in task-level constructs while maintaining a taxonomy-agnostic aggregation rule, the framework is inherently adaptable: new benchmarks, modalities, or task definitions can be incorporated without modifying the underlying notion of general intelligence.